OpenStudy (anonymous):

Help with partial fractions integrals? I keep getting weird coefficients...(problem below)

OpenStudy (anonymous):

The site keeps kicking me off... but anyway my original problem was \[\int\limits_{}^{}\frac{ 10 }{ (x-1)(x^2+9) }dx\] and I set up the partial fractions like this \[10=A(x^2+9)+B(x-1)\] but the coefficients don't add up, I get A and B both equalling 0 yet somehow 9A-B equalling 10. How do I do this?

OpenStudy (phi):

I think you set it up like this \[ \frac{A}{x-1} + \frac{Bx + C}{x^2 +9} \] putting these over the common denominator (x-1)(x^2+9) we get \[ A(x^2+9) + (Bx+C)(x-1) \] it looks like you left out the C expanding, and collecting powers of x, and equating to 10 \[ Ax^2 +9A +Bx^2 -Bx + Cx -C = 10 \\ (A+B) x^2 + (C-B)x + (9A-C) = 10 \] from which we get A+B=0, C-B=0 , 9A-C = 10

OpenStudy (anonymous):

oh because its a quadratic on the bottom, ok thank you :)